# Vib2Mol: from vibrational spectra to molecular structures—a versatile deep learning model

### v3

@@ Appendix A Details about Datasets
VB-QM9: 133,434 organic small molecules extracted from QM9[33], composed of C, H, O, N, and F atoms, with the number of heavy atoms less than 10. Each molecule in this subset has only one stable conformation
VB-ZINC15: 50,114 drug molecules extracted from ZINC15[34], involving a wider range of elements, including C, H, O, N, S, F, Cl, Br, P, and Si, with the number of heavy atoms ranging from 4 to 45. Notably, since the ZINC15 dataset contains many isomers, and VB-zinc15 only ensures the uniqueness of ZINC-IDs, 7,556 molecules in this subset have multiple stable conformations.
@@
VB-geometry: 6835 organic small molecules extracted from GEOM[35], each with two stable conformations. We randomly used the spectrum of one conformation as the query input and the other as the reference spectrum, thus constructing a test set for evaluating the model’s spectrum-to-spectrum matching performance.
SDBS: 2815 organic molecules extracted from Spectral Database for Organic Compounds (SDBS[17]), which contains experimental Raman and infrared spectra simultaneously. This data was collected up to July 1, 2025.
NIST-IR: 12144 experimental infrared spectrum-molecule pairs extracted from NIST Chemistry WebBook[18]. This data was collected up to July 1, 2025.
@@ Appendix B Ablation study of MLM
Appendix C Details about multi-modal spectral input
The Vib2Mol architecture is designed to handle multimodal spectral data flexibly. As shown in Figure S7, the spectral input is structured as a $B×C×L$ tensor, where: $B$ is the batch size, $C$ is the number of spectral channels (modalities), $L$ is the spectral dimension (the number of sampling points, which is 1024 in this work). We use a 1D convolution with an equal kernel size and stride to convert the input tensor into patch tokens. This convolutional layer’s input channel count is $C$ and its output channel count is $D$ (the feature dimension of each token, which is 768 in our case).
For a single-modality input, $C=1$. For a multimodal input, $C=2$, as the Raman ($S_{r}aman$) and infrared ($S_{I}R$) spectra are concatenated along the channel dimension ($S_{i}n=concat(S_{r}aman,S_{I}R)$). The key reason we can simply adjust the value of $C$ without other architectural changes is that both Raman and infrared spectra share the same x-axis (wavenumber).
@@ Appendix D Figures
Figure S12: Diagram of processing multimodal spectral input.
Appendix E Tables
Table S3: Benchmarking on QM9S
@@ Appendix E Tables
Table S4: Benchmarking on VB-Mols
@@
Table S5: Benchmarking on VB-GEOM
extra info Spectrum-spectrum de novo
retrieval generation
Raman IR Raman IR
Cosine similarity None 36.62 34.28 - -
Pearson correlation coefficient None 35.86 33.39 - -
VibraCLIP None 74.96 64.43 - -
PBSA chemical formula - - 50.9 47.32
IR2Mol chemical formula - - 0.10 0.07
SMEN None 41.80 49.01 0.00 0.00
Vib2Mol None 77.54 75.33 60.23 56.14
Vib2Mol chemical formula 77.54 75.33 66.86 63.07
Table S6: Benchmarking on SDBS
@@
Table S7: Benchmarking on NIST-IR

### v4

@@ Appendix A Details about Datasets
VB-QM9: 133,434 organic small molecules extracted from QM9[ramakrishnan2014quantum], composed of C, H, O, N, and F atoms, with the number of heavy atoms less than 10. Each molecule in this subset has only one stable conformation
VB-ZINC15: 50,114 drug molecules extracted from ZINC15[sterling2015zinc], involving a wider range of elements, including C, H, O, N, S, F, Cl, Br, P, and Si, with the number of heavy atoms ranging from 4 to 45. Notably, since the ZINC15 dataset contains many isomers, and VB-zinc15 only ensures the uniqueness of ZINC-IDs, 7,556 molecules in this subset have multiple stable conformations.
@@
VB-geometry: 6835 organic small molecules extracted from GEOM[axelrod2022geom], each with two stable conformations. We randomly used the spectrum of one conformation as the query input and the other as the reference spectrum, thus constructing a test set for evaluating the model’s spectrum-to-spectrum matching performance.
SDBS: 2815 organic molecules extracted from Spectral Database for Organic Compounds (SDBS[saito2011development]), which contains experimental Raman and infrared spectra simultaneously. This data was collected up to July 1, 2025.
NIST-IR: 12144 experimental infrared spectrum-molecule pairs extracted from NIST Chemistry WebBook[linstorm1998nist]. This data was collected up to July 1, 2025.
@@ Appendix B Ablation study of MLM
Appendix C SMILES augmentation
We utilized SMILES randomization (also known as SMILES enumeration) as a data augmentation strategy. One molecular structure can be represented by multiple equivalent SMILES strings. For example, propanal can be denoted as “CCC=O”, “C(=O)CC”, or “C(CC)=O”. We randomly sample these equivalent representations during training. This approach encourages the model to learn structural invariance, a technique previously shown in models like IR2Mol and PBSA to significantly enhance generative performance. Detailed implementation scripts are provided here.
def smiles_augment(smiles, augmentation_ratio=0.5):
if random.random() < augmentation_ratio:
mol = Chem.MolFromSmiles(smiles)
if mol:
# Generate randomized Kekule SMILES
augments = Chem.MolToSmiles(mol,
isomericSmiles=False,
canonical=False,
kekuleSmiles=True,
doRandom=True)
return augments
return smiles
@@ Appendix D Figures
Figure S12: Experimental Raman spectra of peptides.
Table S4: Add caption
VibraCLIP SMEN Vib2Mol
Molecular representation atomic numbers and bond features atomic numbers and coordinates SMILES
(2D graph) (3D graph)
Molecular encoder DimeNet++ EGNN Transformer
Spectral encoder MLP Transformer Transformer
num of molecular layers 4 5 6
num of spectral layers 1 5 6
Dimension of molecular embedding 128 256 768
Dimension of spectral embedding 1262 512 768
Dropout rate 0 0.1 0.1
Optimizer AdamW AdamW AdamW
Learning rate 0.0004 0.0001 0.0001
Batch size 128 200 1024
Patch size / spectral length 1 / 1750 7 / 1024 8 / 1024
spectral range 400 4000 400 4000 400 4000
Augment types None None None
Table S5: Hyper-parameters about generation models.
PBSA IR2Mol Vib2Mol
num of heads 8 8 8
num of layers 4 6 6
Embedding dimension 512 512 768
FFN dimension 2048 2048 2048
Dropout rate 0.1 0.1 0.1
Optimizer Adam AdamW AdamW
Learning rate 0.0001 0.001 0.0001
Batch size 128 128 128
Patch size / spectral length 64 / 3200 75 / 1625 8 / 1024
spectral range 400 3980 650 3900 400 4000
Augment types SMILES SMILES SMILES
Table S6: Spectrum-structure retrieval and de novo generation performance on the QM9S test set, reported as Top-1 Recall (%).
@@
Table S7: Spectrum-structure retrieval and de novo generation performance on the VB-Mols test set, reported as Top-1 Recall (%).
@@
Table S8: Spectrum-spectrum retrieval performance on the VB-GEOM test set, reported as Top-1 Recall (%).
Spectrum-spectrum
retrieval
Raman IR
Cosine similarity 36.62 34.28
Pearson correlation coefficient 35.86 33.39
VibraCLIP 75.22 70.16
SMEN 41.8 49.01
Vib2Mol 77.54 75.33
Table S9: Spectrum-structure retrieval and de novo generation performance on the SDBS test set, reported as Top-1 Recall (%)
@@
Table S10: Spectrum-structure retrieval and de novo generation performance on the NIST-IR test set, reported as Top-1 Recall (%).
@@
Table S11: Large-scale retrieval performance of Vib2Mol-MM against the PubChem database.
Additional molecules from PubChem Recall@1 / Recall@5
0 (only original test set) 95.74 / 97.87
1,000 95.39 / 96.45
10,000 89.36 / 90.43
100,000 83.69 / 84.40
1,000,000 68.44 / 69.86
10,000,000 48.23 / 51.42
80,368,968 (total PubChem) 30.60 / 33.81
Table S12: Latent space similarity analysis of experimental peptides. Values represent mean cosine similarity (%) between a sample’s embedding and the rest of the dataset.
Structural similarity Spectral similarity
Dipeptide 55.74±19.91 50.34±21.23
Tripeptide 51.69±24.04 49.45±22.92
Tetrapeptide 51.51±24.04 49.10±22.92
Averaged 53.64±22.19 49.86±22.27
@@
Table S14: Comparison of model size, computational complexity, and memory requirements.
Training phase Computational cost (GFLOPS) # of parameters (M) # of active parameters (M) Memory requirements (G/GPU)
CL-only alignment 13.2 66.66 66.66 36.87
CL+Matching alignment 35.02 114.1 114.1 75.27
Vib2Mol alignment 35.02 145.86 81.03 69.03
LM-only generation 15.95 80.48 80.48 5.42
MLM-only generation 16.04 80.48 80.48 5.36
LM+MLM generation 23.16 128.36 128.36 7.96
Vib2Mol generation 27.47 145.86 113.55 9.08